OpenStudy (anonymous):

Solve:

OpenStudy (anonymous):

11/8x=36

OpenStudy (anonymous):

btw its one and one eighth

OpenStudy (anonymous):

is the x part of the denominator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok, so 1 1/8 as a mixed number is 9/8

OpenStudy (anonymous):

so that means the problem can be written as 9/8 x = 36. If you multiply both sides by the reciprocal of 9/8 you get x = 36 (8/9) or x = 32.

OpenStudy (anonymous):

so the answer is 32

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thx can u help me with the rest

OpenStudy (anonymous):

I'm going to be on and off, but if you post your questions in a new post I'm sure others will help out as well.

OpenStudy (anonymous):

ok